OpenStudy (anonymous):

Write an exponential function that goes through the point (0,4).

OpenStudy (anonymous):

\[y = x ^{2} + 4\]

OpenStudy (anonymous):

Wouldn't an exponential function be in the form of y=ab^x? The x should be power?

OpenStudy (anonymous):

Oops I meant y=b^x

OpenStudy (anonymous):

Oh, I see what you mean: \[y = a(b)^{x}\] In that case: \[f(0) = 4 = a(b)^{0}\] \[4 = a(1)\] \[a = 4\] So no we get: \[f(x) = 4(b)^{x}\] And 'b' can be anything you like, it doesn't really matter in this case because when it's exponentiated to 0, we always get 1.

OpenStudy (anonymous):

Correction: *now

OpenStudy (anonymous):

Oh, I understand now. Thank you very much!